# Generalization of Task Parameterized Dynamical Systems using Gaussian Process Transportation

### v1

@@ IV Policy Transportation
where $\bm{\gamma}$ and $\bm{\psi}$ are, respectively, a linear and nonlinear transformation. The fitting of the function is made in two steps: first the linear transformation $\bm{\gamma}(\bm{x})$ is obtained, and then the nonlinear transformation $\bm{\psi}(\bm{\gamma}(\bm{x}))$ is fitted on the residual error.
IV-A Linear Transformation
@@
$\underbrace{\bm{\mathcal{T}}-\bm{\bar{\mathcal{T}}}}_{y_{\text{label}}}=\bm{\gamma}(\underbrace{\bm{\mathcal{S}}-\bar{\bm{\mathcal{S}}}}_{x_{\text{label}}})$ (8)
@@
$\bm{U},\bm{\Sigma},\bm{V}^{\top}=(\bm{\mathcal{S}}-\bm{\bar{\mathcal{S}}})^{\top}(\bm{\mathcal{T}}-\bm{\bar{\mathcal{T}}})$ (9)
@@
$\bm{A}=\bm{V}\bm{U}^{\top},$ (10)
however, if $\det(\bm{A})<0$, the last column of $\bm{V}$ is flipped in sign, and the computation of the rotation matrix is repeated. This ensures the transformation is a proper rotation matrix without any reflection; see [29] for more details. Hence, the linear transformation on any point in the space can be computed as
@@
Fig. 3,c shows a linear transformation of the source and a grid of points from the original space depicted in Fig. 3,b.
IV-B Non-linear Transportation
@@
$\underbrace{\bm{\mathcal{T}}-\bm{\gamma}(\bm{\mathcal{S}})}_{y_{\text{label}}}=\bm{\psi}(\underbrace{\bm{\gamma}(\bm{\mathcal{S}})}_{x_{\text{label}}}).$ (12)
The nonlinear function $\bm{\psi}$ can be any nonlinear regressor, such as a Neural Network, a Random Forest, a Gaussian Process, etc. However, the inducting bias given by the nature of the nonlinear function will affect the regression output when going out of distribution, i.e., far away from the given data. For example, suppose the function is approximate with a GP with a distance-based kernel $k$, such as a square exponential kernel.
If the prior distribution is set to be a zero-mean function, when making predictions, see (2) in regions of the space far away from the source distribution points, the final transportation converges to just being a linear transformation, see Fig. 3,d. Knowing the out-of-distribution (o.o.d.) properties of our transportation policy is desirable, considering that we will transport points that are not necessarily close to the point of the source/target distribution.
IV-C Transportation of the Dynamics
Although the transportation map allows the transport of any point of the original demonstration in the new situation, for example, to generalize the demo on cleaning a new surface, we still have not formulated a transportation function for the velocity field $\bm{\dot{x}}$. It is not as trivial as computing the numerical differentiation of the transported trajectories. We consider the policy labels as independent points, no longer part of a trajectory. This allows us to learn from multiple demonstrations and to change the velocity label connected to them if providing (teleoperated) feedback or aggregating new data from interactive demonstrations [25].
Nevertheless, the partial derivative of the transportation mapping can be exploited in the velocity field generalization.
$\mathcal{\phi}$${f}$${\hat{f}}$$\frac{\partial\mathcal{\phi}}{\partial x}$$x$$\dot{x}$$\hat{x}$$\dot{\hat{x}}$
Figure 4: Mathematical Scheme of Policy Transportation.
Figure 5: Graphical representation of the mathematical scheme of policy transportation.
Figure 6: Standard Deviation quantification on the velocity field. Transportation Uncertainty was computed with Eq. (16) and quantifies the (heteroscedastic) uncertainty on the transported label (velocity) corresponding to the transported demonstration. Epistemic Uncertainty is the resulting model uncertainty when fitting the new policy $\hat{f}$. Total Uncertainty is the resulting standard deviation after computing the variance sum of transportation and epistemic uncertainties.
Given the transportation function defined in the source space and projecting in the target space, i.e.,
$\bm{\hat{x}}=\bm{\phi(x)}$
by differentiating w.r.t. time on both sides and using the chain rule, we obtain the velocity field in the transported space as
$\bm{\dot{\hat{x}}}=\frac{\partial\bm{\phi}(\bm{x})}{\partial\bm{x}}\bm{\dot{x}}=\bm{J}(\bm{x})\bm{\dot{x}}$ (13)
@@ IV-C Transportation of the Dynamics
Fig. 6 summarizes all the steps done during transportation that can also be visualized in Fig. 6 on a 2D example. In the first row of the two figures, the original demonstration is used to compute the vector field. Then, in the second row, the demonstration and the vector field are transported, thanks to the transportation map $\bm{\phi}$ fitted on the observed source and target distribution.
In particular, after the transportation of $\hat{\bm{x}}$ and $\dot{\hat{\bm{x}}}$, in the vertical lines of scheme in Fig. 6, the new policy, denoted as $\bm{\hat{f}}$, is fitted again using the transported $\bm{\hat{x}}$ and $\bm{\dot{\hat{x}}}$ labels.
IV-D Robot Orientation and Stiffness Generalization
However, when learning and controlling the Cartesian robot pose, we must also generalize the desired end-effector orientation.
@@ IV-D Robot Orientation and Stiffness Generalization
Let us consider the end effector to be a vector of infinitesimal length with the base $\bm{x_{0}}$ on the end effector position and pointing in the direction of the robot orientation during the demonstration, $\bm{R}_{ee}$. The transportation of the tip of the vector can be obtained using the Taylor approximation of Eq. (7), according to
$\bm{\hat{x}_{\text{tip}}}=\bm{\phi(x_{\text{tip}})}\approx\bm{\phi(x_{0})}+\frac{\partial\bm{\phi}}{\partial\bm{x}}{\bm{\epsilon}}=\bm{\phi(x_{0})}+\bm{J}\bm{R}_{ee}{\bm{\epsilon}_{0}}$ (14)
@@
where $\bm{R}_{ee}\in\mathbb{R}^{3\times 3}$ is the original robot orientation, ${\bm{\epsilon}_{0}}$ is a vector with infinitesimal dimension that has zero orientation. From Eq. (14), it is readily apparent that the transported orientation matrix of the robot end-effector becomes
$\bm{\hat{R}}_{ee}:=\bm{J}\bm{{R}}_{ee}.$ (15)
The transported orientation matrix needs to be orthogonal with the determinant equal to 1; hence, the pre-multiplication matrix $\bm{J}$ must have the same properties. We enforce this by normalizing $\bm{J}$ with its determinant and finding the corresponding orthogonal matrix with a QR decomposition.
@@
Additionally, when implementing policies on a Cartesian impedance control, the stiffness $\bm{K}_{\text{s}}$ and the damping matrix $\bm{{D}}$ must also be transported. The change of coordinates of the stiffness and the damping follows from the transportation of the robot-applied force on the environment found using Hooke’s law, i.e.,
$\bm{\hat{F}}_{s}=\bm{\hat{K}}_{s}\Delta\bm{\hat{x}}=\bm{\hat{K}}_{s}\overbrace{\bm{J}\Delta\bm{x}}^{\Delta\bm{\hat{x}}}=\bm{J}\overbrace{\bm{K}_{s}\Delta\bm{x}}^{\bm{F}_{s}}.$
@@
$\bm{\hat{K}_{\text{s}}}=\bm{J}{\bm{K}_{\text{s}}}\bm{J}^{T}$
@@
IV-E Transportation Uncertainty
A probabilistic function approximator, like a GP, will also provide the uncertainty on transportation output from Eq. (3) that can be propagated in the transported dynamical system. In particular, a GP derivative is also a GP [26] and its existence will depend on the differentiability of the kernel function. The correlation between derivative samples can be expressed as the second partial derivative $k^{11}=\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}k(x_{i},x_{j})$ while the correlation between derivative samples and function samples is $k^{10}=\frac{\partial}{\partial x_{i}}k(x_{i},x_{j})$. Thus, the mean and variance prediction of the derivative of the Gaussian Process become
@@ IV-E Transportation Uncertainty
The uncertainty quantification of the transportation policy becomes essential for calculating the final uncertainty on the control variable, e.g., the velocity. In Fig. 6, the uncertainty is also displayed as a shaded area around the demonstration and as the “warmness” of the color in the vector field.
The uncertainty of the velocity labels is due to the propagation of the original labels through the derivative of the (uncertain) transportation map of (13), i.e.,
$\bm{\Sigma}_{\hat{\bm{x}}}=\bm{\Sigma}_{\frac{\partial\bm{\phi}(\bm{x})}{\partial\bm{x}}}\bm{\dot{x}^{2}}.$ (17)
given the definition of the weighted sum of Gaussian variables [30].
Hence, considering that the labels are uncertain, the prediction of the resulting heteroscedastic GP [31] can be computed as the sum of the epistemic and (variable) aleatoric uncertainty, that is
Fig. 6 depicts the transportation uncertainty on the norm of the velocity, calculated with Eq. (16) and the epistemic uncertainty of the model $\bm{\hat{f}}$, computed with Eq. (3) using transported position and transported velocities labels. From Fig. 6 and 6, it is possible to appreciate that the transportation uncertainties grow when evaluating in regions that are far away from the task parameterization points since the transportation is less certain when going far away from the distribution data; on the other hand, the epistemic uncertainty grows when evaluating in points that are far from the transported demonstration.
V 2-D Simulations and comparisons
The availability of (calibrated) uncertainties is an important feature to improve the trustworthiness in deploying robot motion generalization. In this section, we evaluate Gaussian Process Transportation (GPT) on generalizing the demonstration in a 2D surface cleaning task and on a reference frame-to-frame motion generalization. Our goal for these simulated experiments is
  • to illustrate and compare how the generalization process differs when employing regressors other than a Gaussian Process or methodologies from the state-of-the-art while generalizing a cyclic demonstration that approaches and then retreats from the surface “to clean”, in Sec. V-A;
  • assess and compare GPT’s ability to generalize in multi-reference frame tasks, measuring its performance against state-of-the-art algorithms, in Sec. V-B.
V-A 2-D surface cleaning
Fig. 6 visualizes the transportation of the given demonstration, in red, from the source to the target space, using the transportation map $\bm{\phi}$ where the non-linear component was chosen to be a GP, given the out-of-distribution prediction and the calibrated uncertainty quantification. However, other state-of-the-art function approximators can be used to fit the transportation function without loss of generality. To ensure a fair comparison, the mean linear transformation, i.e., $\bm{\gamma}$, is applied to all trajectories before using the different methods to perform the non-linear transportation. Table I summarises the method with their properties, while Fig. 7 shows the generalization of the demonstration when these methods are used.
Kernelized Movement Primitives (KMP) [6], in this study, fits the motion as a function of time while Laplacian editing (LE) [13] considers the topology of the demonstration to be a chain, i.e., a graph where only consecutive vertices are connected with an edge or as a ring, when the demonstration is periodic, i.e. also starting and ending nodes are connected, like in Fig. 7. Hence, every point of the source distribution is matched with the closest point of the demonstration. Then, each point of the trajectory, or the graph, is moved, knowing the new desired target location of the matched points.
Hence, LE and KMP do not provide any uncertainty on the transportation process.
All the other transportation regressors, a part of the GP, are ensembles of popular regression functions, i.e., Ensemble Random Forest (E-RF). Ensemble Neural Networks (E-NN) and Ensemble Neural Flows (E-NF). An ensemble is a collection of multiple individual models, trained independently, whose combined predictions are used to estimate a distribution on the prediction, i.e., mean and variance. In this example, Neural Networks are simple multi-layer perceptrons while Neural Flows [32] are bijective neural networks, i.e., flows, usually used to learn a mapping from a simple probability distribution to a more complex target distribution.
Fig. 7 depicts the mean and the uncertainty bounds of 2-$\sigma$ for the transported trajectories when using ensembles and GPs. The bounds are computed from the different fitted models in the ensembles while it is computed analytically for the GP and the depicted GP samples of Fig. 7 are drawn from the posterior distribution.
@@
From Fig. 7, the reader can appreciate how the GP is the only regressor with well-calibrated and unbiased epistemic uncertainty quantification and minimal mean prediction distortion of the trajectory when transporting points far away from the source distribution. For example, the E-NN, has higher uncertainty on the right side of the demonstration, even though the points are at the same distance from the surface, while E-RF generates an un-distorted overconfident transformation, i.e., the uncertainty does not grow when going out of distribution.
Figure 7: Qualitative comparison of transportation of demonstration in target space for 2D surface cleaning. The colored lines are the samples of the final transported trajectory policy, i.e. $\hat{x}=\phi(x)$, and the orange area is 2 standard deviation. The black curve is the 1-D surface to clean.
Method Modality Vel. Gen. Transportation Uncertainty
KMP [6] way-points
LE [13] way-points
E-RF [33] continuous estimated
E-NF [24] continuous estimated
E-NN [34] continuous estimated
GP [26] continuous analytical
V-B Multiple Reference Frames
In the literature, one of the main applications of task parameterization is the generalization w.r.t. one or more reference frames. For example, if we teach a robot how to pour water into a glass, we want the robot to automatically generalize the motion w.r.t. any glass position. The task, in this particular case, can be parameterized with the location of the reference frames of each object, which is necessary to track for a successful generalization of the motion. Typically, the motion is projected in any of the reference frames, and a policy is learned w.r.t. each of the frames, leaving out the decision on the relevance of each frame for every timestep. Task-Parametrized Gaussian Mixture Model (TP-GMM) learns a GMM model for the projected demonstration for each of the frames and, during executions, the Gaussians of each frame are combined using the property that the product of Gaussian (PoG) is still a Gaussian, see [2] for more details. Given the GMM, different control formulations are possible, for example only relying on the current state of the system, i.e., $\bm{\Delta}\bm{x}_{i}=\bm{f}(\bm{x}_{i})$ [35] [2] or by using a Hidden Markov Model (HMM) formulation that also considers the progress during the execution of the trajectory $\bm{x}_{i+1}=\bm{f}(\bm{x}_{i},\bm{\alpha}_{i})$, where $\bm{\alpha_{i}}$, in the context of a mixture mode, selects the properties of the model (mean and variance) that can be used in a tracking algorithm, such as a Linear Quadratic Regulator (LQR) [2, 36]. However, the latent transition matrix between the different states of the HMM is unknown. They need to be estimated using a forward pass algorithm, i.e., the Viterbi algorithm [36], that requires an initial guess trajectory to infer the most likely state transition that generated that initial guess and again generate the most likely motion according to the model. However, having an initial guess can be prohibitive when evaluating the movement in a novel configuration of starting and goal frames.
Differently from these task-parameterized approaches, our proposed method does not track only the reference frame but a set of points that are relevant to the starting and goal object. To guarantee a fair comparison with the state of the art, in Fig. 8, when generalizing using GPT, only 5 points are tracked w.r.t. each reference frame, capturing the position but also the local orientation of the frames. In Fig. 8, what we describe as DMP uses the same mathematical structure of GPT but only relies on a linear transformation, which is why the result is not able to capture the non-linear deformation due to the frame orientation. One of the main perks of our proposed method is the ability to generalize any dynamical system generated by even only one demonstration, unlike GMM-based methods where, to capture a meaningful mixture model, at least two diverse demonstrations need to be provided. Additionally, the GMM uncertainty of the final multi-frame model that results from the PoG does capture the uncertainty of the transportation, while, as depicted in Fig. 8, the GP transportation results in growing uncertainties when transporting points of the demonstrations that are less correlated with the source-target points.
Fig. 8 highlights the discrepancy in the performance of GMM methods on the training set and the test set. At the same time, the reproduction of a known combination of the frames results in accurate rollouts of the policies both when executing them as a dynamical system (TP-GMM) [^1] Implementation available at https://github.com/BatyaGG/Task-Parameterized-Gaussian-Mixture-Model [35] then as an optimal tracking problem of a multi-transition Hidden Markov Model (HMM) [^2] Implementation available at https://gitlab.idiap.ch/rli/pbdlib-python/-/blob/master/notebooks/ [36], when evaluating on the test set, generated on random reorganization of the frames, the resulting trajectories do not successfully reach the goal frame neither in position or orientation. To quantify and compare the different methods, we conducted a quantitative analysis comparing the generalization on known trajectories from the demonstration set or the reaching performance on a randomly generated frame set.

### v2

@@ III-E Learning Transportation Maps: $\phi$
The inference of the function is made in two steps: first, the affine transformation $\bm{\gamma}(\bm{x})$ is obtained, and then the nonlinear transformation $\bm{\psi}(\bm{\gamma}(\bm{x}))$ is fitted only on the residual error.
Affine transformation
Figure 2: 2D transportation. Distribution match depicts the source and the target distribution correspondence used to train the transportation function. Source Distribution depicts a grid of points in the original space. Linear Transformation shows the effect on the original grid when only an affine transformation $\gamma$ is used to match source and target distribution. Nonlinear Transportation captures the deformation of the space when the source points are forced to match the target ones via the application of the complete policy $\phi$.
@@
$\bm{\gamma}:\mathbb{R}^{3}\to\mathbb{R}^{3}\text{ s.t. }\bm{\mathcal{T}}-\bm{\bar{\mathcal{T}}}=\bm{\gamma}(\bm{\mathcal{S}}-\bar{\bm{\mathcal{S}}}).$ (4)
@@
$\bm{U}\bm{\Sigma}\bm{V}^{\top}=(\bm{\mathcal{S}}-\bm{\bar{\mathcal{S}}})^{\top}(\bm{\mathcal{T}}-\bm{\bar{\mathcal{T}}})$ (5)
@@
$\bm{A}=\bm{V}\bm{U}^{\top}.$ (6)
However, if $\det(\bm{A})<0$, the last column of $\bm{V}$ is flipped in sign, and the computation of the rotation matrix is repeated. This ensures that the transformation is a proper rotation matrix without any reflection; see [46] for more details. Hence, the linear transformation on any point in the space can be computed as
@@
However, if $\det(\bm{\Sigma})=0$, i.e., $\bm{\Sigma}$ is not full rank, it implies that the rotation is not uniquely defined and is set to the identity by default. Fig. 2,c shows a linear transformation of the source and a grid of points from the original space depicted in Fig. 2,b.
Having a global affine transform reduces the necessary complexity of the nonlinear function. In fact, the norm of the distance between the target and the roto-translated source is lower or equal to the norm of the distance between the target and the original source, i.e.,
$||\bm{\mathcal{T}}-\bm{\gamma}(\bm{\mathcal{S}})||\leq||\bm{\mathcal{T}}-\bm{\mathcal{S}}||.$
The proof is a direct consequence that the affine transformation transforms the original source to be as close as possible to the target, given the linear constraint.
Nonlinear transformation
@@
$\bm{\psi}:\mathbb{R}^{3}\to\mathbb{R}^{3}\text{ s.t. }\bm{\mathcal{T}}-\bm{\gamma}(\bm{\mathcal{S}})=\bm{\psi}(\bm{\gamma}(\bm{\mathcal{S}})).$ (8)
The nonlinear function $\bm{\psi}$ can be any nonlinear regressor, such as a Multi-layer Perceptron (MLP), a Gaussian Process (GP) or a Locally Weighted Translation (LWT) [31]. However, the inductive bias given by the nature of the nonlinear function will affect the regression output when transporting points far away from the source distribution. For example, suppose that the function is approximated with a GP with a distance-based kernel $k$, such as a square exponential kernel (see App. A). If the prior distribution is set to be a zero-mean function when making predictions in regions of the space far away from the source distribution points, the final transportation converges to just being an affine transformation, see Fig. 2,d. Knowing the out-of-distribution (o.o.d.) properties of our policy transportation is desirable, considering that we will transport points of the policy that are not necessarily close to the point of the source/target distribution. If no keypoints are provided, the zero deformation prior is applied to the space.
III-F Transportation of Labels: Position and Velocity
This paper focuses on generalizing the policy labels so that a new policy can be fitted to satisfy the new circumstances. Given our set of position labels $\bm{\mathcal{X}}$ from the original policy, the transported labels $\bm{\mathcal{\hat{X}}}$ are then obtained as
$\bm{{\mathcal{\hat{X}}}}=\bm{\phi(\mathcal{X})}.$
Although the transportation map allows the transport of any point of the original demonstration in the new situation, for example, to generalize the demo on cleaning a new surface, we still have not formulated a transportation function for the velocity labels $\bm{\mathcal{\dot{X}}}$. This is not as trivial as computing the numerical differentiation of the transported trajectories. We consider the policy labels as independent points, no longer part of a trajectory. This allows us to learn from multiple demonstrations and to change the velocity label through feedback or aggregating new data from interactive demonstrations [1] [4]. Nevertheless, the partial derivative of the transportation mapping can be exploited in the velocity field generalization.
Given the transportation function defined in the source space and projecting in the target space and by differentiating w.r.t. time on both sides and using the chain rule, we obtain the velocity labels in the transported space as
$\bm{\mathcal{\dot{\hat{X}}}}=\frac{\partial\bm{\phi}(\bm{\mathcal{X}})}{\partial\bm{\mathcal{X}}}\bm{\mathcal{\dot{X}}}=\bm{J}(\bm{\mathcal{X}})\bm{\mathcal{\dot{X}}}$ (9)
@@ III-F Transportation of Labels: Position and Velocity
Fig. 5 and Fig. 5 summarize how the transportation function $\bm{\phi}$ is used to transport the label of the original policy. We assume to have a set of labels used to fit the original dynamical system, i.e. $\bm{\dot{x}}=f(\bm{x})$; we show how to transport these labels using the transportation function $\bm{\phi}(\bm{x})$ before fitting the new dynamical system $g(\bm{x})$. If the original policy labels were sampled from a stable dynamical system, a sanity check on transportation is to ensure that the determinant of the Jacobian is larger than zero for any of the transported labels [34]; this also means that the function is locally invertible, hence locally diffeomorphic. If the map is diffeomorphic and the original labels were sampled from a stable dynamic system, they will maintain the stability property. However, given the new set of labels, if the task requires the motion to be asymptotically stable, then the proper regression [10], [8], [9], [31] can be selected to enforce this requirement (even if the whole labels do not preserve this property). When using a function approximator that allows setting a zero-prior and evaluating out of distribution, the Jacobian converges to the average rotation matrix, i.e., $\bm{J}\approx\bm{A}$. Fig. 5 shows the vector fields obtained by fitting $f$ and $g$ using a vanilla GP but different regressors can be chosen, like LPV [10] or NN [8].
III-G Transportation of Labels: Orientation, Stiffness, and Damping
$\mathcal{\phi}$${f}$$g$$\frac{\partial\mathcal{\phi}}{\partial x}$$\bm{\mathcal{X}}$$\bm{\mathcal{\dot{X}}}$$\bm{\mathcal{\hat{X}}}$$\bm{\mathcal{\dot{\hat{X}}}}$
Figure 3: Mathematical Scheme of Policy Transportation.
Figure 4: Graphical representation of the mathematical scheme of policy transportation. The color temperature in the fields represents the uncertainty of the prediction.
Figure 5: Standard Deviation quantification on the velocity field. Transportation Uncertainty and quantifies the (heteroscedastic) uncertainty on the transported label (velocity) corresponding to the transported demonstration. Epistemic Uncertainty is the resulting model uncertainty when fitting the new policy $\hat{f}$. Total Uncertainty is the resulting standard deviation after computing the variance sum of transportation and epistemic uncertainties.
@@ III-G Transportation of Labels: Orientation, Stiffness, and Damping
When learning and controlling the Cartesian robot pose, we must also generalize the desired end-effector orientation. Let us consider the end effector to be a vector of infinitesimal length with the base $\bm{x_{0}}$ on the end effector position and pointing in the direction of the robot orientation during the demonstration, $\bm{R}$. The transportation of the tip of the vector can be obtained using the Taylor approximation of Eq. (2), according to
$\bm{\hat{x}_{\text{tip}}}=\bm{\phi(x_{\text{tip}})}\approx\bm{\phi(x_{0})}+\frac{\partial\bm{\phi}}{\partial\bm{x}}{\bm{\epsilon}}=\bm{\phi(x_{0})}+\bm{J}\bm{R}{\bm{\epsilon}_{0}}$ (10)
@@
where $\bm{R}\in\mathbb{R}^{3\times 3}$ is the original robot orientation, ${\bm{\epsilon}_{0}}$ is a vector with infinitesimal dimension that has zero orientation.
[…]
From Eq. (10), it is readily apparent that the transported orientation labels of the robot end-effector become
$\bm{\mathcal{\hat{R}}}=\bm{J_{\perp}}\bm{\mathcal{{R}}}.$ (11)
@@
Additionally, when implementing policies on a Cartesian impedance control, the stiffness $\bm{K}$ and the damping matrix $\bm{{D}}$ must also be transported. The change of coordinates of the stiffness and the damping follows from the transportation of the robot-applied force on the environment, found using Hooke’s law, i.e.,
$\bm{\hat{F}}_{s}=\bm{\hat{K}}\Delta\bm{\hat{x}}=\bm{\hat{K}}\overbrace{\bm{J}\Delta\bm{x}}^{\Delta\bm{\hat{x}}}=\bm{J}\overbrace{\bm{K}\Delta\bm{x}}^{\bm{F}_{s}}.$
@@
$\bm{\hat{\mathcal{K}}}=\bm{J_{\perp}}{\bm{\mathcal{K}}}\bm{J_{\perp}}^{T}$
@@
III-H Transporting Uncertainty
Given the reduced set of available keypoints, some parts of the space that are not close to them will be transported with a certain degree of uncertainty. A probabilistic model, like a Gaussian Process (GP) or any approximator of its posterior like an Ensemble Neural Network (E-NN) [47], will also provide the uncertainty on transported labels that are used to fit the new motion policy. In particular, a GP derivative is also a GP [48] and its existence will depend on the differentiability of the kernel function, see Appendix A. On the other hand, the uncertainty of the derivative in an ensemble neural network (E-NN) is computed as the standard deviation of the derivatives of each model.
The uncertainty quantification of the transported labels becomes essential for calculating the final uncertainty on the control variable, e.g., the velocity. In Fig. 5, the uncertainty is also displayed as a shaded area around the demonstration and as the “warmness” of the color in the vector field. The uncertainty of the velocity labels is due to the propagation of the original velocity labels through the derivative of the (uncertain) transportation map of Eq. (9), i.e.,
$\bm{\Sigma}_{\hat{\bm{x}}}=\bm{\Sigma}_{\frac{\partial\bm{\phi}(\bm{x})}{\partial\bm{x}}}\bm{\mathcal{\dot{X}}}^{2}.$ (12)
given the definition of the weighted sum of Gaussian variables [49]. Hence, considering that the labels are uncertain, the prediction of the resulting policy can be computed as the sum of the epistemic (due to the reduced set of policy labels) and aleatoric uncertainty (due to the reduced set of keypoints), that is
Fig. 5 depicts the transportation uncertainty on the norm of the velocity and the epistemic uncertainty of the model $\bm{\hat{f}}$ using transported position and transported velocities labels. From Fig. 5 and 5, it is possible to appreciate that the transportation uncertainties grow when evaluating in regions that are far away from the task parameterization points since the transportation is less certain when going far away from the distribution data; on the other hand, the epistemic uncertainty grows when evaluating in points that are far from the transported demonstration.
IV 2-D Simulations and comparisons
The availability of (calibrated) uncertainties is an important feature that improves trustworthiness in deploying robot motion generalization. In this section, we evaluate different maps, for example, defining a Gaussian Process Transportation (GPT) for generalizing the demonstration in a 2D surface cleaning task and on a multiple reference motion generalization. Our goal for these simulated experiments is
  • to illustrate and compare how the generalization process differs when employing regressors other than a Gaussian Process or methodologies from the state-of-the-art while generalizing a cyclic demonstration that approaches and then retreats from the surface “to clean”, in Sec. IV-A;
  • assess and compare the policy transportation ability to generalize in multi-reference frame tasks, measuring its performance against state-of-the-art algorithms, in Sec. IV-B.
IV-A 2-D surface cleaning
Figure 6: Qualitative comparison of the transportation of a demonstration in target space for 2D surface cleaning. The colored lines are the samples of the final transported trajectory policy, i.e. $\hat{x}=\phi(x)$, and the orange area is two standard deviations. The black curve is the 1-D surface to clean.
Method Modality Vel. Transport Diffeomorphic Uncertainty
Reshaped-KMP linear transform + trajectory-reshaping analytical
LE [16] linear transform + trajectory-reshaping
LTW [31] [33] linear transform + bijective map
E-Flow [50, 32] linear transform + bijective map estimated
E-NN [51] linear transform + residual map estimated
GPR transform + residual map analytical
Fig. 5 visualizes the transportation of the given demonstration, in red, from the source to the target space, using the transportation map $\bm{\phi}$ where the non-linear component was chosen to be a GP, given the out-of-distribution prediction and the calibrated uncertainty quantification. However, other state-of-the-art function approximators can be used to fit the transportation function without loss of generality. To ensure a fair comparison, the mean linear transformation, i.e., $\bm{\gamma}$, is applied to all trajectories before using the different methods to perform the non-linear transportation. Table I summarises the method with their properties, while Fig. 6 shows the generalization of the demonstration when these methods are used. The illustrated trajectories are the result of the transportation of the original demonstration.
The figure is organized in this way:
  • the first column shows the result of trajectory-based reshaping methods, Kernelized Movement Primitives (KMP) and Laplacian Editing (LE);
  • the second column shows the result by using diffeomorphic maps;
  • the third column shows the results by using generic regression models, i.e., NN, and GPs.
KMP [7], in this study, fits the motion as a function of time, i.e. $\mathcal{X}=f(t)$, while Laplacian editing (LE) [17, 16] considers the topology of the demonstration to be a chain, i.e., a graph where only consecutive vertices are connected with an edge or as a ring, when the demonstration is periodic, i.e. also starting and ending nodes are connected, like in Fig. 6. For both of these trajectory-based methods, every point of the source distribution is matched with the closest point of the demonstration, solving the Hungarian assignment problem. Then, each point of the trajectory, or the graph, is moved, knowing the new desired target location of the matched points by using Laplacian editing or the time-based equivalent for the transportation process, which we define as Reshaped-KMP, and it is governed by these equations:
$\hat{\mathcal{X}}=\mathcal{X}+\phi(\bm{t}),\quad\text{where}\quad\phi(\bm{t_{*}})=\mathcal{T}-\mathcal{S}$
where $\bm{t}_{*}$ corresponds to the time indexes in the trajectory of the selected elements to modify, and $\bm{t}$ is the complete set of time indexes of the recorded trajectory.
Figure 7: Comparisons using many more surfaces. The first two rows used trajectory reshaping methods, while the last two used the proposed transportation method with a GP and an LWT. We also report the computational time [s] to fit and transform the regressor, the accuracy in matching overlapping source points with target points, and when using transportation maps, the percentage of transported points with a Jacobian larger than zero.
@@ III-H Transporting Uncertainty
From Fig. 6, the reader can appreciate how the GP-based methods are the only regressor with well-calibrated and unbiased epistemic uncertainty quantification and minimal mean distortion of the trajectory when transporting points far away from the source distribution. For example, the E-NN, has higher uncertainty on the right side of the trajectory, even though the points are at the same distance from the surface. The Neural Flows show an undesired distortion of the demonstration when evaluating far from the surface. Differently, the LWT performs an undistorted generalization but does not provide any measure of uncertainty.